# A Survey on Intent-aware Recommender Systems

### v1

@@ 3.1. Profile Diversity Matching
In terms of real-world applications, little is known in the literature about how large-scale platforms like YouTube technically diversify their recommendations to cater for time-varying user intents. Various real-world online services, e.g., Netflix or Amazon, however apparently address this problem by providing multiple recommendation lists to users at a time (Gomez-Uribe and Hunt, 2015). At least some of these lists are clearly connected to a particular assumed user intent, e.g., “continue watching” or “discover our new arrivals”. Even though such multi-list approaches are common in practice, only a few works on the topic can be found in the academic literature, e.g., (Jannach et al., 2021a; Starke et al., 2021a; Rahdari et al., 2022).
@@ 3.2. Sequence- and Time-Aware Approaches
A number of intent-aware approaches can also be found in the area of sequential recommendation settings. Like in the previously discussed session-based scenarios, sequential recommendation approaches are based on time-ordered user interaction logs and the general goal is to predict the user’s immediate next action. However, these approaches do not assume that the data is organized in sessions. The differentiation of long-term preferences and short-term intents is central to several works also in sequential recommendation settings. Zhang et al. (Zhang et al., 2019), for instance, propose a corresponding architecture that comprises two components to capture a user’s transient interests based on their observed interaction trajectories: a self-attention model for short-term intents and a latent factor model for long-term preferences. Long-term and short-term aspects are also combined by Zhang et al. (Zhang et al., 2018) and by Liu et al. (Liu et al., 2021), where the latter develop a novel approach to consider time intervals between observed events. Other approaches involve sophisticated temporal reasoning approaches as well as model the evolution of user behavior over time. These models may additionally also rely on item meta-data to factor in semantic relationships between items (Wang et al., 2021c) or time-varying popularity of items (Jiang et al., 2021). The use of item meta-data, in particular in the form of category information, is central to a number of other works (Guo et al., 2020; Chen et al., 2022a, 2019; Xu et al., 2021a; Cai et al., 2021). In many of these works, the time-varying interest in different item categories is used as a proxy for user intent, and the goal is thus to model item category trajectories in the interaction data. Technically, graph-based approaches and attention are commonly used. Moreover, different types of side information are taken into account, including social information (Chen et al., 2022a), purchase cycles (Guo et al., 2020) or information extracted from tags or user reviews (Li et al., 2022d). Some works also consider multiple possible category-related intents in parallel (Xu et al., 2021a) and are able to process multiple types of user actions (e.g., clicks, purchases etc.)(Chen et al., 2019; Xu et al., 2021a). Finally, a few proposals exist that consider domain-specific aspects. Zhu et al. (Zhu et al., 2023a), for example, focus on Point-of-Interest recommendation problems and their model incorporates a “locational intent” signals from spatio-temporal information and location categories. Ding et al. (Ding et al., 2022), on the other hand, concentrate on the fashion domain. They consider shopping-specific intents (e.g., “match” or “substitute”) and furthermore incorporate automatically extracted features of fashion items, which are used to model item sequences on a level not tied to individual items. Lastly, two intent-aware approaches for predicting the next app usage on mobile phones are presented in (Lee and Lee, 2014) and (Changmai et al., 2019). In (Lee and Lee, 2014), a system based on (sequential) rules is proposed that predicts the next user action, i.e., app invocation, based on the user’s predicted intent and context. A related approach is presented in (Changmai et al., 2019), which relies both on spatio-temporal and sequential information to predict the next user intent.
@@ 3.3. Latent Intent Modeling
The DisenHAN model by Wang et al. (Wang et al., 2020d) features a particular disentangled graph attention network that leverages different aspects in a Heterogenous Information Network (HIN). The output of their layered architecture are a number of intent-enhanced embedding vectors, which represent different aspects of a user and items. A related approach to intent-aware recommendation based on knowledge graphs was proposed by (Wang et al., 2021b). The work is conceptually similar to the DGCF approach discussed above (Wang et al., 2020c), but was designed to also considers fine-grained item information in the intent-modeling process. Knowledge graphs and intent-awareness were also in the focus of Zhang et al. (Zhang et al., 2022b, a). In their models, the user and item representations are disentangled to different spaces, resulting in a number of intent embeddings, which are then used to obtain enhanced embeddings that are finally used for recommendation. Similar ideas were later adopted also by Li et al. (Li et al., 2023b), who create intent graphs for each user before training based on connected entities in the knowledge graph. A KG-based method inspired by topic modeling was proposed by Li et al. (Li et al., 2023c), which aims to overcome potential problems when extracting intent information from graph relations, as done, e.g., in (Wang et al., 2021b).
@@
Various alternative technical approaches to intent-aware sequential recommendation are proposed in (Cen et al., 2020; Wang et al., 2022c; Chen et al., 2023; Ma et al., 2020; Chang et al., 2023). Cen et al. (Cen et al., 2020) focus on the candidate generation phase of recommender system. They explore two methods for multi-interest extraction, one based on self-attention, as done in several other works, and one in which the multiple interests of a given user are viewed as capsules (Sabour et al., 2017). In this latter approach, the dynamic routing method from CapsNet is used to generate multiple output capsules that correspond to interest embeddings. Wang et al. (Wang et al., 2022c) later challenge the effectiveness of the greedy inference method used, for example, in (Cen et al., 2020) and propose a target-interest distillation approach to dynamically aggregate multi-interest embeddings for a given context. Li et al. (Chen et al., 2023) instead present another work based on the attention mechanism, and they specifically use multi-head attention, where each head corresponds to one of a limited set of attentions to users. Ma et al. (Ma et al., 2020), like  (Cen et al., 2020), focus on candidate generation, and they challenge the effectiveness of multi-head attention in models such as SASRec (Kang and McAuley, 2018) for intent modeling. Instead, they propose a specific intention disentanglement layer appended after single-head SASRec encoder, which involves an intention clustering and weighting step for a given small set of latent categories. In yet another technical approach, Chang et al. (Chang et al., 2023) put forward a probabilistic approach in which user intent is modeled as latent variables, which connect past observed user behavior and future behavior. Finally, Oh et al. (Oh et al., 2022) propose to model what they call “implicit session contexts” in session-aware recommendation settings. An implicit session context is seen to be equivalent to a latent intent, and the model uses a next-context predictor to guide the next-item prediction model.
@@ 4.1. Domains, Datasets, and Model Inputs
Figure 2 shows the application focus of current research in IARS. We note that one paper can be assigned to more than one application domain. The analysis shows that e-commerce scenarios are the main focus and driver of research in this area. This distinguishes the field of IARS apart from other surveys, where movie recommendations were traditionally the most common application domain (Klimashevskaia et al., 2023; Jannach et al., 2012). We attribute the focus on e-commerce domain mainly to the increased interest in session-based and sequential recommendation problems, which are prevalent in e-commerce environments. In the residual Others category, we find individual works that for example focus on specific problems such as next-app prediction or ad recommendation.
@@
Except for those few cases mentioned above, where users are expected to explicitly specify their current intent, a given user’s intent is assumed to be unknown and must be estimated from the observed user behavior, from contextual factors or from different forms of side information. Considering the surveyed literature comprising 81 papers, the following observations can be made.
In more than 30 papers, i.e., in above one third of the cases, no information beyond the observed interaction signals (e.g., item views, purchases, reviews) and their temporal ordering or timestamps is taken into account. These cases, for instance, include works on session-based and sequential recommendation, where potential user intents are modeled through item co-occurrences in sessions or where the goal is to guess the main intent of a given session, for example through an attention mechanism (Li et al., 2017; Liu et al., 2018; Oh et al., 2022). Also several approaches that aim at modeling multiple current interests of users, which create multiple intent-based embeddings, or which propose specific network layers for intent modeling, do not leverage any side information, e.g. (Wang et al., 2022d; Zhu et al., 2022). In terms of the used interaction signals, most works only consider one type of interaction, e.g., item views, even though datasets like YOOCHOOSE or Diginetica exist, which comprise multiple types of user interactions. Interestingly, among all surveyed papers, only a handful consider more than one type of interaction in parallel, e.g., (Mehrotra et al., 2019; Yang et al., 2021; Xu et al., 2021b; Loyola et al., 2017; Shi and Ghedira, 2017; Liu et al., 2021).
Leveraging additional information about users and items beyond interaction signals is however not uncommon in IARS research. About half of the papers incorporate some form of item meta-data. Most commonly, the category of the items is a central feature, and in several works the observed interest in items of a certain category is used as a proxy for the user’s unknown intent. Other types of item meta-data, e.g., the genre of a movie, are commonly used in knowledge graph-based approaches to intent-aware recommendation. Other types of information are only considered in a small set of works, like user demographics (Li et al., 2019; Liu et al., 2021), item content/concepts (Li et al., 2022d; Wei et al., 2022; Hu et al., 2020), user tags and reviews (Kaya and Bridge, 2018; Wu et al., 2022; Guo et al., 2020; Volokhin and Agichtein, 2018a), social information (Li et al., 2021; Chen et al., 2022a; Picot-Clémente et al., 2014b; Changmai et al., 2019; Li et al., 2022a), context (location and time) (Lee and Lee, 2014; Li et al., 2022b; Picot-Clémente et al., 2014b; Zhu et al., 2023b), or item popularity (Jiang et al., 2021).
@@ 4.2. Experimental Approaches
Finally, in addition to the work by Shi et al. (Shi and Ghedira, 2017) discussed above, survey data was collected in  (Mehrotra et al., 2019) and (Bénédict et al., 2023b). The work by Mehrotra et al. (Mehrotra et al., 2019) reports on findings of a study at Spotify. A sample of three million Spotify users were presented with a survey in the app, where they were asked (a) about their satisfaction with their experience on the home screen, and (b) about the reasons of using the app on this day. For the latter question, six predefined options were presented which were identified through interviews with 12 Spotify users. Over 100.000 users responded to the survey, providing insightful information about the distribution of user intents in a real-world environment. Furthermore, the survey helped to identify user intents that were not considered in the preceding interview-based research. The entire research by Mehrotra et al. was reproduced later in the video domain by Benedict et al. (Bénédict et al., 2023b). For their study, they also first interviewed user experience experts about possible user intents in this domain, and then conducted an in-app survey to obtain the distribution of intents and to learn about additional intents they had not thought of before.
@@ 4.4. Reproducibility
Considering other potentially problematic research practices, we found that in many surveyed papers that provide results of offline experiments, nothing is mentioned about the tuning process of the ‘state-of-the-art’ baselines. Furthermore, in several works, the same embedding sizes are used across compared models, even though embedding sizes are a hyperparameter that has to be tuned for each model and dataset (Shehzad and Jannach, 2023). [^17] We deliberately refrain from highlighting individual papers here. The list of studied papers can be found online. Overall, these observations may lead to worries about the true progress that is achieved in the field of IARS in the last years. The research community therefore seems to be in need of in-depth reproducibility studies in this area.
@@ 5.1. Discussion and Gaps in Current Research
An apparent research gap, as discussed above, lies in the fact that today’s research on intent-aware recommender systems is done almost entirely through offline experiments, often with limited reproducibility. Moreover, in these offline experiments we observed a strong focus on a single quality dimension, i.e., recommendation accuracy. It is therefore highly important that the research community more often conducts experiments with humans in the loop to truly understand the value of intent-awareness for users in different dimensions. On a positive note, our survey surfaced a few works which were validated in field tests with deployed systems. Unfortunately, several of the described A/B tests lasted only a relatively short amount of time, e.g., two or three weeks. Also, in several cases, central details such as the target KPI are not reported.
@@
In terms of application domains, the research scope on IARS is not different from the general literature on recommender systems, with a focus on the media and e-commerce domains. We also found that much of the literature is also based on well-known and widely used datasets. As these datasets do not contain explicit information about user intents, the user’s intent has to be derived from the available information about users, items and user-item interactions. A certain gap therefore exists in terms of richer datasets, which either contain explicit information about intent [^18] Explicit intent information could be acquired through in-app surveys, as in (Mehrotra et al., 2019) or additional interaction signals that can be used for intent prediction. Interestingly, current research often does not leverage all the pieces of information that can be found in existing datasets, as mentioned above. Only a few works consider multiple types of behaviors in their models, e.g., clicks, purchases, or add-to-cart actions in e-commerce settings. Furthermore, some datasets like Diginetica [^19] https://competitions.codalab.org/competitions/11161 contain rich information like user search queries. Intuitively, in particular search queries represent an important predictor for user intent, but current research does not yet exploit such information to a large extent.

### v3

@@ 3.1. Profile Diversity Matching
In terms of real-world applications, little is published in the literature on how large-scale platforms like YouTube technically diversify their recommendations to cater for time-varying user intents. Various real-world online services, e.g., Netflix or Amazon, however apparently address this problem by providing multiple recommendation lists to users at a time (Gomez-Uribe and Hunt, 2015). At least some of these lists are clearly connected to a particular assumed user intent, e.g., “continue watching” or “discover our new arrivals”. Even though such multi-list approaches are common in practice, only a few works on the topic can be found in the academic literature, e.g., (Jannach et al., 2021a; Starke et al., 2021a; Rahdari et al., 2022).
@@ 3.2. Sequence- and Time-Aware Approaches
A number of intent-aware approaches can also be found in the area of sequential recommendation settings. Like in the previously discussed session-based scenarios, sequential recommendation approaches are based on time-ordered user interaction logs and the general goal is to predict the user’s immediate next action. However, these approaches do not assume that the data is organized in sessions. The differentiation of long-term preferences and short-term intents is central to several works also in sequential recommendation settings. Zhang et al. (Zhang et al., 2019), for instance, propose a corresponding architecture that comprises two components to capture a user’s transient interests based on their observed interaction trajectories: a self-attention model for short-term intents and a latent factor model for long-term preferences. Long-term and short-term aspects are also combined by Zhang et al. (Zhang et al., 2018) and by Liu et al. (Liu et al., 2021), where the latter develop a novel approach to consider time intervals between observed events. Other approaches involve sophisticated temporal reasoning approaches as well as model the evolution of user behavior over time. These models may additionally also rely on item meta-data to factor in semantic relationships between items (Wang et al., 2021c) or time-varying popularity of items (Jiang et al., 2021). The use of item meta-data, in particular in the form of category information, is central to a number of other works (Guo et al., 2020; Chen et al., 2022a, 2019; Xu et al., 2021b; Cai et al., 2021). In many of these works, the time-varying interest in different item categories is used as a proxy for user intent, and the goal is thus to model item category trajectories in the interaction data. Technically, graph-based approaches and attention are commonly used. Moreover, different types of side information are taken into account, including social information (Chen et al., 2022a), purchase cycles (Guo et al., 2020) or information extracted from tags or user reviews (Li et al., 2022d). Some works also consider multiple possible category-related intents in parallel (Xu et al., 2021b) and are able to process multiple types of user actions (e.g., clicks, purchases etc.)(Chen et al., 2019; Xu et al., 2021b). Finally, a few proposals exist that consider domain-specific aspects. Zhu et al. (Zhu et al., 2023a), for example, focus on Point-of-Interest recommendation problems and their model incorporates a “locational intent” signals from spatio-temporal information and location categories. Ding et al. (Ding et al., 2022), on the other hand, concentrate on the fashion domain. They consider shopping-specific intents (e.g., “match” or “substitute”) and furthermore incorporate automatically extracted features of fashion items, which are used to model item sequences on a level not tied to individual items. Lastly, two intent-aware approaches for predicting the next app usage on mobile phones are presented in (Lee and Lee, 2014) and (Changmai et al., 2019). In (Lee and Lee, 2014), a system based on (sequential) rules is proposed that predicts the next user action, i.e., app invocation, based on the user’s predicted intent and context. A related approach is presented in (Changmai et al., 2019), which relies both on spatio-temporal and sequential information to predict the next user intent.
@@ 3.3. Latent Intent Modeling
The DisenHAN model by Wang et al. (Wang et al., 2020d) features a particular disentangled graph attention network that leverages different aspects in a Heterogenous Information Network (HIN). The output of their layered architecture are a number of intent-enhanced embedding vectors, which represent different aspects of a user and items. A related approach to intent-aware recommendation based on knowledge graphs was proposed by (Wang et al., 2021b). The work is conceptually similar to the DGCF approach discussed above (Wang et al., 2020c), but was designed to also consider fine-grained item information in the intent-modeling process. Knowledge graphs and intent-awareness were also in the focus of Zhang et al. (Zhang et al., 2022b, a). In their models, the user and item representations are disentangled to different spaces, resulting in a number of intent embeddings, which are then used to obtain enhanced embeddings that are finally used for recommendation. Similar ideas were later adopted also by Li et al. (Li et al., 2023b), who create intent graphs for each user before training based on connected entities in the knowledge graph. A KG-based method inspired by topic modeling was proposed by Li et al. (Li et al., 2023c), which aims to overcome potential problems when extracting intent information from graph relations, as done, e.g., in (Wang et al., 2021b). Finally, Lin et al. (Lin et al., 2023) propose an intent-aware approach that re-ranks and refines a recommendation list by learning latent user intent representations from text reviews. Their work is unique in that it uses user-generated content in the form of reviews as auxiliary information.
@@
Various alternative technical approaches to intent-aware sequential recommendation are proposed in (Bhattacharya et al., 2017; Cen et al., 2020; Sulikowski et al., 2021; Wang et al., 2022c; Chen et al., 2023; Ma et al., 2020; Chang et al., 2023). Cen et al. (Cen et al., 2020) focus on the candidate generation phase of recommender system. They explore two methods for multi-interest extraction, one based on self-attention, as done in several other works, and one in which the multiple interests of a given user are viewed as capsules (Sabour et al., 2017). In this latter approach, the dynamic routing method from CapsNet is used to generate multiple output capsules that correspond to interest embeddings. Wang et al. (Wang et al., 2022c) later challenge the effectiveness of the greedy inference method used, for example, in (Cen et al., 2020) and propose a target-interest distillation approach to dynamically aggregate multi-interest embeddings for a given context. Li et al. (Chen et al., 2023) instead present another work based on the attention mechanism, and they specifically use multi-head attention, where each head corresponds to one of a limited set of attentions to users. Ma et al. (Ma et al., 2020), like  (Cen et al., 2020), focus on candidate generation, and they challenge the effectiveness of multi-head attention in models such as SASRec (Kang and McAuley, 2018) for intent modeling. Instead, they propose a specific intention disentanglement layer appended after single-head SASRec encoder, which involves an intention clustering and weighting step for a given small set of latent categories. In yet another technical approach, Chang et al. (Chang et al., 2023) put forward a probabilistic approach in which user intent is modeled as latent variables, which connect past observed user behavior and future behavior. Oh et al. (Oh et al., 2022), on the other hand, propose to model what they call “implicit session contexts” in session-aware recommendation settings. An implicit session context is seen to be equivalent to a latent intent, and the model uses a next-context predictor to guide the next-item prediction model.
The works by Bhattacharya et al. (Bhattacharya et al., 2017) and Sulikowski et al. (Sulikowski et al., 2021) stand out in that they use very specific ways of considering the users’ past and ongoing interaction behavior for intent-modeling and prediction. Bhattacharya et al. focus on a particular use case, where textual reports are recommended to users. Their intent modeling approach considers both navigational aspects, e.g., which reports users have interacted with and for how long, as well as content-related aspects. A particularity of their use case is that the authors equate the item to be recommended, i.e., a specific report, with the predicted intent or user goal. This aspect is similar to the work by Sulikowski et al., who aim to predict the intent of users to purchase individual items, which is similar to the classical recommendation problem. However, the purchase probabilities in their work are based on an intent-modeling approach that considers fine-grained user interaction signals such as clicks, dwelling times, mouse moves, and scrolling activities. A further unique aspect of their approach also lies in the use of a fuzzy rule-learning technique.
@@ 4.1. Domains, Datasets, and Model Inputs
Figure 2 shows the application focus of current research in IARS. We note that one paper can be assigned to more than one application domain. The analysis shows that e-commerce scenarios are the main focus and driver of research in this area. This distinguishes the field of IARS from other surveys, where movie recommendations were traditionally the most common application domain (Klimashevskaia et al., 2023; Jannach et al., 2012). We attribute the focus on e-commerce domain mainly to the increased interest in session-based and sequential recommendation problems, which are prevalent in e-commerce environments. Furthermore, the availability of relevant datasets and widely-accepted baseline models may have contributed to the observed focus on e-commerce as an application area. In the residual Others category, we find individual works that for example focus on specific problems such as next-app prediction or ad recommendation.
@@
Interestingly, while many of the surveyed works address problems that are highly relevant in practice—in particular sequential recommendation in e-commerce—research in the area of IARS is still largely driven by academia. Only a handful of papers are authored or co-authored by research groups from industry. Correspondingly, we also did find only a number of papers where IARS models were evaluated in an A/B test in a production environment, as we will discuss later.
@@
Except for those few cases mentioned above, where users are expected to explicitly specify their current intent, a given user’s intent is assumed to be unknown and must be estimated from the observed user behavior, from contextual factors or from different forms of side information. Considering the surveyed literature comprising 85 papers, the following observations can be made.
In more than 30 papers, i.e., in above one third of the cases, no information beyond the observed interaction signals (e.g., item views, purchases, reviews) and their temporal ordering or timestamps is taken into account. These cases, for instance, include works on session-based and sequential recommendation, where potential user intents are modeled through item co-occurrences in sessions or where the goal is to guess the main intent of a given session, for example through an attention mechanism (Li et al., 2017; Liu et al., 2018; Oh et al., 2022). Also several approaches that aim at modeling multiple current interests of users, which create multiple intent-based embeddings, or which propose specific network layers for intent modeling, do not leverage any side information, e.g. (Wang et al., 2022c; Zhu et al., 2022). In terms of the used interaction signals, most works only consider one type of interaction, e.g., item views, even though datasets like YOOCHOOSE or Diginetica exist, which comprise multiple types of user interactions. Interestingly, among all surveyed papers, only a handful consider more than one type of interaction in parallel, e.g., (Mehrotra et al., 2019; Yang et al., 2021; Xu et al., 2021b; Loyola et al., 2017; Shi and Ghedira, 2017; Liu et al., 2021).
Leveraging additional information about users and items beyond interaction signals is however not uncommon in IARS research. About half of the papers incorporate some form of item meta-data. Most commonly, the category of the items is a central feature, and in several works the observed interest in items of a certain category is used as a proxy for the user’s unknown intent. Other types of item meta-data, e.g., the genre of a movie, are commonly used in knowledge graph-based approaches to intent-aware recommendation. Other types of information are only considered in a small set of works, like user demographics (Li et al., 2019; Liu et al., 2021), item content/concepts (Li et al., 2022d; Wei et al., 2022; Hu et al., 2020), user tags and reviews (Kaya and Bridge, 2018; Wu et al., 2022; Guo et al., 2020; Volokhin and Agichtein, 2018a), social information (Li et al., 2021; Chen et al., 2022a; Picot-Clémente et al., 2014a; Changmai et al., 2019; Li et al., 2022a), context (location and time) (Lee and Lee, 2014; Li et al., 2022b; Picot-Clémente et al., 2014a; Zhu et al., 2023b), or item popularity (Jiang et al., 2021).
@@ 4.2. Experimental Approaches
Finally, in addition to the work by Shi et al. (Shi and Ghedira, 2017) discussed above, survey data was collected in  (Mehrotra et al., 2019) and (Bénédict et al., 2023a). The work by Mehrotra et al. (Mehrotra et al., 2019) reports on findings of a study at Spotify. A sample of three million Spotify users were presented with a survey in the app, where they were asked (a) about their satisfaction with their experience on the home screen, and (b) about the reasons of using the app on this day. For the latter question, six predefined options were presented which were identified through interviews with 12 Spotify users. Over 100.000 users responded to the survey, providing insightful information about the distribution of user intents in a real-world environment. Furthermore, the survey helped to identify user intents that were not considered in the preceding interview-based research. The entire research by Mehrotra et al. was reproduced later in the video domain by Benedict et al. (Bénédict et al., 2023a). For their study, they also first interviewed user experience experts about possible user intents in this domain, and then conducted an in-app survey to obtain the distribution of intents and to learn about additional intents they had not thought of before.
@@ 4.4. Reproducibility
Considering other potentially problematic research practices, we found that in many surveyed papers that provide results of offline experiments, nothing is mentioned about the tuning process of the ‘state-of-the-art’ baselines. Furthermore, in several works, the same embedding sizes are used across compared models, even though embedding sizes are a hyperparameter that has to be tuned for each model and dataset (Shehzad and Jannach, 2023). [^16] We deliberately refrain from highlighting individual papers here. Overall, these observations may lead to worries about the true progress that is achieved in the field of IARS in the last years. The research community therefore seems to be in need of in-depth reproducibility studies in this area.
@@ 5.1. Discussion and Gaps in Current Research
An apparent research gap, as discussed above, lies in the fact that today’s research on intent-aware recommender systems is done almost entirely through offline experiments, often with limited reproducibility. Moreover, in these offline experiments we observed a strong focus on a single quality dimension, i.e., recommendation accuracy. It is therefore highly important that the research community more often conducts experiments with humans in the loop to truly understand the value of intent-awareness for users in different dimensions. On a positive note, our survey surfaced at least a few works which were validated in field tests with deployed systems. Unfortunately, several of the described A/B tests lasted only for a relatively short period of time, such as two or three weeks. Also, in several cases, central details such as the target KPI are not reported. We note that running an A/B test for three weeks may be considered sufficient in some settings, in particular when a large enough user base is involved in the experiment. However, papers often do not mention if there has been a ramp-up phase when a new model is deployed and do not report how the KPIs develop over time. If such factors are not considered, the long-term effects of a new model may easily be overestimated. For example, we may observe more interest (clicks) immediately after a new model is deployed, when users explore the new types of recommendations. However, the more interesting and often unanswered question is if the new model truly has a lasting effect on user behavior. Generally, the limited number of field tests in the literature call for more intensive collaborations between academia and industry, both to evaluate new IARS models in a production environment and to create new real-world datasets for academic research.
@@
In terms of application domains, the research scope on IARS is not different from the general literature on recommender systems, with a focus on the media and e-commerce domains. We also found that much of the literature is also based on well-known and widely used datasets. As these datasets do not contain explicit information about user intents, the user’s intent has to be derived from the available information about users, items and user-item interactions. A certain gap therefore exists in terms of richer datasets, which either contain explicit information about intent [^17] Explicit intent information could be acquired through in-app surveys, as in (Mehrotra et al., 2019) or additional interaction signals that can be used for intent prediction. Interestingly, current research often does not leverage all the pieces of information that can be found in existing datasets, as mentioned above. Only a few works consider multiple types of behaviors in their models, e.g., clicks, purchases, or add-to-cart actions in e-commerce settings. Furthermore, some datasets like Diginetica [^18] https://competitions.codalab.org/competitions/11161 contain rich information like user search queries. Intuitively, in particular search queries represent an important predictor for user intent, but current research does not yet exploit such information to a large extent. Another possible source of information for intent modeling and prediction is to rely on fine-grained logs of user interactions, such as clicks, mouse movements, or scrolling behavior. Such interaction logs were for example used in (Bhattacharya et al., 2017) and (Sulikowski et al., 2021). Finally, Sun et al. (Lin et al., 2023) demonstrated a novel way to leverage external auxiliary information for intent modeling, in their case in the form of textual reviews.
Generally, both for already existing and future datasets, it is crucial to validate that these datasets are suited for the given task. In the area of sequential recommendation models, it was recently pointed out that much of the research done in the past few years is based on datasets that may actually not be truly sequential in nature (Woolridge et al., 2021; Hidasi and Czapp, 2023; Klenitskiy et al., 2024). Similar considerations apply for datasets used in IARS research, which often requires sequential data.
@@
On a more general level, a long-term ambition in algorithm-oriented research in recommender systems—and in other areas of applied machine learning—is to predict which of several models would perform best, given only the characteristics of the available data. Some work in that direction can for example be found in (Adomavicius and Zhang, 2012), but the problem in general seems largely unresolved. Translated to the domain of IARS, it would be similarly valuable to be able to know which approach—using implicit or explicit intents—should be adopted for a given use case by only considering the types of data and recorded interactions that are available. In the surveyed literature, we found no work that even tried to empirically compare the performance of explicit and implicit intent modeling approaches.